# Robust Signal Detection with Quadratically Convex Orthosymmetric Constraints

### v2

@@ Appendix B Deferred Proofs of the Lemmas and Main Theorems
where (i) uses (A.1). Now assuming $\mathbb{P}_{\nu}\ll\mathbb{P}_{0}$, and let $L_{\nu}=\frac{d\mathbb{P}_{\nu}}{d\mathbb{P}_{0}}$ be the Radon-Nikodym derivative of $\mathbb{P}_{\nu}$ to $\mathbb{P}_{0}$. By Cauchy-Schwartz inequality, we have
@@
B.3 Proof of Theorem 3.3
In this section, we prove Theorem 3.3. Our proof relies on Lemma B.1 and a carefully designed optimal coupling. To establish Lemma B.1, we first recall the notion of differential privacy in B.2 and present two auxiliary lemmas, Lemma B.3 and Lemma B.4. Based on these tools, we then construct a symmetric set using the QCO properties and compute the Hamming distance as an application of Lemma A.5. The computation proceeds by constructing a sequence of couplings that reduces the evaluation of the Hamming distance into that of the total variation distance.
Lemma B.1 (main lemma).
@@
$\mathbb{E}_{(\mathbf{X},\mathbf{X}^{\prime})\sim(\mathcal{U},\mathcal{V})}d_{H}(\mathbf{X},\mathbf{X}^{\prime})\lesssim\epsilon N.$ (B.1)
@@
To prove Lemma B.1, we start with the notion of $(\epsilon,\delta)$-differential privacy.
Definition B.2 ($(\epsilon,\delta)$-differential privacy).
For a pair of non-negative real numbers $(\epsilon,\delta)$, a randomized algorithm $\mathcal{A}$ that is defined on the sample space $\Omega$ is a $(\epsilon,\delta)$-differentially private algorithm if for any $S\subset\text{range}(\mathcal{A})$ and any $X,X^{\prime}\in\Omega$ with $d_{H}(X,X^{\prime})\leq 1$, we have
$\mathbb{P}\left(\mathcal{A}(X)\in S\right)\leq e^{\epsilon}\mathbb{P}\left(\mathcal{A}(X^{\prime})\in S\right)+\delta.$ (B.2)
When $\delta=0$, the algorithm is called pure differential privacy; when $\epsilon=0$, the algorithm is called $\delta$-fully approximate differential privacy.
With this concept and motivated by Theorem 11 in Acharya et al., (2018), we are able to prove a lemma that is very close to Lemma B.1.
Lemma B.3.
@@
$\mathbb{E}_{(\mathbf{X},\mathbf{X}^{\prime})\sim(\mathcal{U},\mathcal{V})}d_{H}(\mathbf{X},\mathbf{X}^{\prime})\leq D.$ (B.3)
Then, if a robust test $\mathcal{A}:\Omega\mapsto\left\{0,1\right\}$ between $\mathcal{U}$ and $\mathcal{V}$ is a $\delta$-fully approximate differential private algorithm, with type-I and type-II error probabilities both below $0.1$, then we must have $D=\varOmega(\frac{1}{\delta})$.
@@
$\mathbb{P}\left(\mathcal{A}(\mathbf{X})=0\cap\mathcal{A}(\mathbf{X}^{\prime})=1\cap d_{H}(\mathbf{X},\mathbf{X}^{\prime})<10D\right)\geq 0.7.$ (B.4)
@@
$|\mathbb{P}\left(\mathcal{A}(\mathbf{X})=0\right)-\mathbb{P}\left(\mathcal{A}(\mathbf{X}^{\prime})=0\right)|\leq\delta\cdot d_{H}(\mathbf{X},\mathbf{X}^{\prime}).$ (B.5)
It is worth noting that the probability in (B.4) includes two parts of randomness: the randomness of $\mathbf{X}$ and $\mathbf{X}^{\prime}$ from $\Omega$, and the randomness of $\mathcal{A}$. Now take a realization of $\mathbf{X}$ and $\mathbf{X}^{\prime}$, denoted by $\mathbf{x}$ and $\mathbf{x}^{\prime}$, that satisfy (B.4), we have
$1-0.7\geq\mathbb{P}(\mathcal{A}(\mathbf{x}^{\prime})=0)\geq\mathbb{P}(\mathcal{A}(\mathbf{x})=0)-\delta\cdot d_{H}(\mathbf{x},\mathbf{x}^{\prime})\geq 0.7-\delta\cdot 10D,$ (B.6)
@@
The next lemma shows how to obtain a robust and differentially private test from a given robust test between $\mathcal{U}$ and $\mathcal{V}$ using a black-box robustness-to-privacy transformation. The techniques are from Canonne et al., (2023) and Hopkins et al., (2023).
Lemma B.4.
Suppose $\mathcal{A}$ is a robust test for the mean testing problem (1.2) achieving type-I and type-II errors at most $0.1$ under a contamination level at most $\epsilon$. Then, for $\delta=\frac{1}{\epsilon N}$, there exists a $\delta$-fully approximate differential privacy $\widetilde{\mathcal{A}}$ whose error probabilities are also at most $0.1$ for any dataset with size $N$.
@@
Now assume $\widetilde{\mathbf{X}}$ is from $H_{0}$. By the definition of robust testing, with probability at least $0.9$, we can arbitrarily modify $\widetilde{\mathbf{X}}$ and obtain the same result as long as the number of contamination does not exceed $\epsilon N$. This equivalently says that $\mathcal{S}(\widetilde{\mathbf{X}},\mathcal{A})\geq\epsilon N$, which, by the construction of $\widetilde{\mathcal{A}}$, implies that $\widetilde{\mathcal{A}}$ will output $1$ with probability 0 in this case. On the other hand, if $\widetilde{\mathbf{X}}$ is from $H_{1}$, then with probability at least $0.9$, we will have $\mathcal{A}(X)=1$, in which case $\mathcal{S}(X,\mathcal{A})=0$ and consequently $\widetilde{\mathcal{A}}$ will output $1$ with probability $1$. This shows that $\widetilde{\mathcal{A}}$ is a $\delta$-fully approximate differential privacy with error probabilities at most $0.1$ under both hypotheses. ∎
We are now prepared to prove Lemma B.1.
Proof of Lemma B.1.
Suppose we already have a robust test $\mathcal{A}$ for the mean testing problem 1.2 with the fraction of contamination at most $\epsilon$. Then by Lemma B.4, we can contruct a $\delta$-fully approximate differential privacy $\widetilde{\mathcal{A}}$ with errors at most $0.1$. Therefore, the constructed algorithm $\widetilde{\mathcal{A}}$ satifies the condition of Lemma B.3. Take $D=\mathbb{E}d_{H}(\mathbf{X},\mathbf{X}^{\prime})$ in Lemma B.3, and by the conclusion of Lemma B.3 we should have $\mathbb{E}d_{H}(\mathbf{X},\mathbf{X}^{\prime})\gtrsim\frac{1}{1/\epsilon N}=\epsilon N$. This necessary condition finishes the proof. ∎
@@
$V:=\left\{v\left\lvert\right.v_{i}=\gamma_{i}\theta_{i},\gamma_{i}\in\left\{-1,1\right\},1\leq i\in I\right\},$
where $\theta$ is the vector from Lemma 3.1 setting $c=\frac{(k_{2}^{\star})^{1/4}\sqrt{\epsilon}}{N^{1/4}}\sigma$, i.e., $\left\lVert\theta\right\rVert_{2}=\frac{(k_{2}^{\star})^{1/4}\sqrt{\epsilon}}{N^{1/4}}\sigma$ and $\left\lVert\theta\right\rVert_{\infty}\leq\frac{\sqrt{\epsilon}}{(k_{2}^{\star})^{1/4}N^{1/4}}\sigma$. We are going to compute $d_{H}(\mathbf{X},\mathbf{X}^{\prime})$ using Lemma A.5. Specifically, we compare (in terms of total variation) a distribution of $k$ exchangeable variables generated from $\mathcal{V}$ to a joint distribution of k1k-1 exchangeable variables generated from $\mathcal{N}\left(v,\sigma^{2}\mathbf{I}\right)$ and the last one from $\mathcal{N}\left(0,\sigma^{2}\mathbf{I}\right)$ for $1\leq k\leq N$. We approach this by decomposing both $\mathbf{X}$ and $\mathbf{X}^{\prime}$ by their sequence of conditional probabilities and constructing the optimal coupling between $X_{k}$ and $X^{\prime}_{k}$ conditional on $X_{1},\dots,X_{k-1}$ using Lemma A.5 such that
@@
where $p(x_{k}|x_{1},\dots,x_{k-1})$ is the density of conditional distribution of the first $k$ coordinates of $X^{\prime}$ and $q(x_{k})$ is the density of standard normal distribution. By the definition of total variation distance, it is not hard to see under this optimal coupling that
$\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\sim(\mathcal{U},\mathcal{V})}\mathbf{1}_{\left\{X_{k}\neq X^{\prime}_{k}\right\}}$ $=\mathbb{E}\text{TV}(p(x_{k}|x_{1},\dots,x_{k-1}),q(x_{k}))$
@@
$\mathbb{E}_{(X,X^{\prime})\sim(\mathcal{U},\mathcal{V})}d_{H}(X,X^{\prime})=\mathbb{E}_{X,X^{\prime}\sim(\mathcal{U},\mathcal{V})}\sum\limits_{i=1}^{N}\mathbf{1}_{\left\{X_{i}\neq X^{\prime}_{i}\right\}},$
to evaluate $\mathbb{E}_{(X,X^{\prime})\sim(\mathcal{U},\mathcal{V})}d_{H}(X,X^{\prime})$, it suffices to evaluate
@@
We calculate each individual term. By Pinsker’s inequality A.3, we have
@@
$=\mathbb{E}\ln\left(\frac{\frac{1}{2^{|V|}}\sum\limits_{v\in V}\prod\limits_{i=1}^{k}\exp\left\{-\frac{\left\lVert X_{i}-v\right\rVert_{2}^{2}}{2\sigma^{2}}\right\}}{\frac{1}{2^{|V|}}\sum\limits_{v\in V}\prod\limits_{i=1}^{k-1}\exp\left\{-\frac{\left\lVert X_{i}-v\right\rVert_{2}^{2}}{2\sigma^{2}}\right\}\cdot\exp\left\{-\frac{\left\lVert X_{k}\right\rVert_{2}^{2}}{2\sigma^{2}}\right\}}\right)$
@@
where $(\cdot)_{j}$ denotes the $j$-th coordinate, (i) is from that $\left\lVert v\right\rVert_{2}^{2}$ does not depend on the specfic choice of $v\in V$ by the construction of $V$ and (ii) is from an important observation that
@@
where (i) follows from the Jensen’s inequality. Recall that by Lemma 3.1 we have $\left\lVert\theta\right\rVert_{2}=\frac{(k_{2}^{\star})^{1/4}\sqrt{\epsilon}}{N^{1/4}}\sigma$ and $\left\lVert\theta\right\rVert_{\infty}\leq\frac{\sqrt{\epsilon}}{(k_{2}^{\star})^{1/4}N^{1/4}}\sigma$. Now that $\theta_{i}^{4}\leq\theta_{i}^{2}\cdot\left\lVert\theta\right\rVert_{\infty}^{2}$ for all $i\in\left\{1,2,\dots,d\right\}$, we know $\left\lVert\theta\right\rVert_{4}^{2}\leq\left\lVert\theta\right\rVert_{2}\left\lVert\theta\right\rVert_{\infty}\leq\frac{\epsilon}{\sqrt{N}}\sigma^{2}$. By the definition of $V$, we know that $\left\lVert v\right\rVert_{4}^{2}=\left\lVert\theta\right\rVert_{4}^{2}$. Therefore, we have $\frac{\sqrt{N}\left\lVert v\right\rVert_{4}^{2}}{\sigma^{2}}\leq\frac{\sqrt{N}}{\sigma^{2}}\cdot\frac{\epsilon}{\sqrt{N}}\sigma^{2}\leq\epsilon$. By Lemma B.1, we have
@@
B.4 Proof of Theorem 3.4
In this section, we prove Theorem 3.4. We reach the result by constructing two mixture distributions which are, in fact, identical. We then evaluate the total variation distance between two truncated binomial distributions, which enables us to construct an adaptive adversary, leading high type-II error. Similar techniques also appear in Prasadan and Neykov, (2025) and Chen et al., (2017).
@@
$q_{1}=\frac{(p_{\mu_{0}}-p_{0})\mathbf{1}_{\left\{p_{\mu_{0}}\geq p_{0}\right\}}}{\text{TV}\left(\mathbb{P}_{0},\mathbb{P}_{\mu_{0}}\right)},\ q_{2}=\frac{(p_{0}-p_{\mu_{0}})\mathbf{1}_{\left\{p_{0}>p_{\mu_{0}}\right\}}}{\text{TV}(\mathbb{P}_{0},\mathbb{P}_{\mu_{0}})}.$ (B.7)
@@
$R_{1}:=(1-\epsilon^{\prime\prime})\mathbb{P}_{0}+\epsilon^{\prime\prime}Q_{1},\ R_{2}:=(1-\epsilon^{\prime\prime})\mathbb{P}_{\mu_{0}}+\epsilon^{\prime\prime}Q_{2}.$ (B.8)
It is not hard to verify that actually $R_{1}=R_{2}$, i.e., they represent the same distribution. Consequently $\text{TV}\left(\prod\limits_{i=1}^{N}R_{1}^{i},\prod\limits_{i=1}^{N}R_{2}^{i}\right)=0$, where $\prod\limits_{i=1}^{N}R_{j}^{i},j\in\left\{1,2\right\}$ is the $N$-fold product of $R_{j}$.
@@
Here $B$ is a random variable that follows a binomial distribution with parameters $N$ and $\epsilon^{\prime\prime}$, and (i) follows from the Hoeffding’s inequality.
@@
$0=\text{TV}\left(\prod\limits_{i=1}^{N}R_{1}^{i},\prod\limits_{i=1}^{N}R_{2}^{i}\right)\overset{\text{(ii)}}{\geq}\frac{1}{2}\displaystyle\int_{\mathbb{R}^{d}}\left|S_{1}^{L}-S_{2}^{L}\right|d\mathbf{x}-\frac{1}{2}\displaystyle\int_{\mathbb{R}^{d}}\left|S_{1}^{U}-S_{2}^{U}\right|d\mathbf{x},$
@@
Step IV: conditional argument. Consider the adversary $\mathcal{C}_{B,\epsilon^{\prime\prime}}$ with following strategy. Upon observing the original data $\widetilde{\mathbf{X}}$ and its underlying law (which is either $\mathbb{P}_{0}$ or $\mathbb{P}_{\mu_{0}}$ in our case), the adversary first samples a number $n_{c}$ from the conditional law of $B|B\leq N(\epsilon^{\prime\prime}+1/\sqrt{2N})$, then sample $n_{c}$ replacements from the mixture distribution corresponding to the true underlying law. Specifically, if $\widetilde{\mathbf{X}}$ is from $\mathbb{P}_{0}$, then replacements will be from $Q_{1}$, otherwise the replacements will be from $Q_{2}$. In this way, if we denote $\widetilde{\mathbb{P}}_{0}$ to be the law of the contaminated samples from $\mathbb{P}_{0}$ and $Q_{1}$, and $\widetilde{\mathbb{P}}_{\mu_{0}}$ to be the law of the contaminated samples from $\mathbb{P}_{\mu_{0}}$ and $Q_{2}$, then with minor calculation we have
@@
$\geq 1-\alpha-\text{TV}\left(\widetilde{\mathbb{P}}_{0},\widetilde{\mathbb{P}}_{\mu_{0}}\right)\asymp 1.$
@@
Definition B.5 (Consistent subset).
@@
$E_{e}:=\left\{\left\lVert P^{\star}\sum\limits_{i\in T}X_{i}\right\rVert_{2}^{2}-k^{\star}\left|T\right|\sigma^{2}\geq c\left|T\right|^{2}E^{2}(\epsilon,K,N,\sigma^{2})\right\},$ (B.9)

### v3

@@ Appendix C Deferred Proofs of the Lemmas and Main Theorems
where (i) uses (B.1). Now assuming $\mathbb{P}_{\nu}\ll\mathbb{P}_{0}$, and let $L_{\nu}=\frac{d\mathbb{P}_{\nu}}{d\mathbb{P}_{0}}$ be the Radon–Nikodym derivative of $\mathbb{P}_{\nu}$ to $\mathbb{P}_{0}$. By Cauchy–Schwartz inequality, we have
@@
C.3 Proof of Theorem 3.5
In this section, we formally prove Theorem 3.5. We first recall the core lemma. To establish Lemma C.1, two auxiliary lemmas, Lemma C.3 and Lemma C.4, are presented. Based on these tools, we then construct a symmetric set served as the prior for the mean $\mu$ via the QCO properties and compute the Hamming distance as an application of Lemma B.5 via a sequential optimal coupling. Finally, the lower bound is yielded by the Pinsker’s inequality and a concrete calculation of the KL divergence.
Lemma C.1 (Core lemma).
@@
$\mathbb{E}_{(\mathbf{X},\mathbf{X}^{\prime})\sim(\mathcal{U},\mathcal{V})}d_{H}(\mathbf{X},\mathbf{X}^{\prime})\lesssim\epsilon N.$ (C.1)
@@
We start with a review of the notion of $(\epsilon,\delta)$-differential privacy for convenience.
Definition C.2 ($(\epsilon,\delta)$-differential privacy, Dwork et al., (2006)).
For a pair of non-negative real numbers $(\epsilon,\delta)$, a randomized algorithm $\mathcal{A}$ defined on the sample space $\Omega$ is an $(\epsilon,\delta)$-differentially private algorithm if for any $S\subset\text{range}(\mathcal{A})$ and any $X,X^{\prime}\in\Omega$ with the Hamming distance $d_{H}(X,X^{\prime})\leq 1$, we have
$\mathbb{P}\left(\mathcal{A}(X)\in S\right)\leq e^{\epsilon}\mathbb{P}\left(\mathcal{A}(X^{\prime})\in S\right)+\delta.$ (C.2)
When $\delta=0$, the notion is called pure differential privacy; when $\delta>0$, it is referred to as approximate differential privacy.
Lemma C.3 is motivated by Theorem 11 in Acharya et al., (2018), and shares a nearly identical conclusion with the core lemma.
Lemma C.3.
@@
$\mathbb{E}_{(\mathbf{X},\mathbf{X}^{\prime})\sim(\mathcal{U},\mathcal{V})}d_{H}(\mathbf{X},\mathbf{X}^{\prime})\leq D.$ (C.3)
Then, if a test $\mathcal{A}$ to distinguish between $\mathcal{U}$ and $\mathcal{V}$: $\Omega\mapsto\left\{0,1\right\}$ is a $(0,\delta)$-differentially private algorithm, with type-I and type-II error probabilities both below $0.1$, then we must have $D=\varOmega(\frac{1}{\delta})$.
@@
$\mathbb{P}\left(\mathcal{A}(\mathbf{X})=0\cap\mathcal{A}(\mathbf{X}^{\prime})=1\cap d_{H}(\mathbf{X},\mathbf{X}^{\prime})<10D\right)\geq 0.7.$ (C.4)
@@
$|\mathbb{P}\left(\mathcal{A}(\mathbf{X})=0\right)-\mathbb{P}\left(\mathcal{A}(\mathbf{X}^{\prime})=0\right)|\leq\delta\cdot d_{H}(\mathbf{X},\mathbf{X}^{\prime}).$ (C.5)
It is worth noting that the probability in (C.4) includes two parts of randomness: the randomness of $\mathbf{X}$ and $\mathbf{X}^{\prime}$ from $\Omega$, and the randomness of $\mathcal{A}$. Now take a realization of $\mathbf{X}$ and $\mathbf{X}^{\prime}$, denoted by $\mathbf{x}$ and $\mathbf{x}^{\prime}$, that satisfy (C.4), we have
$1-0.7\geq\mathbb{P}(\mathcal{A}(\mathbf{x}^{\prime})=0)\geq\mathbb{P}(\mathcal{A}(\mathbf{x})=0)-\delta\cdot d_{H}(\mathbf{x},\mathbf{x}^{\prime})\geq 0.7-\delta\cdot 10D,$ (C.6)
@@
Lemma C.4 shows how to obtain a $(0,\delta)$-differentially private test from a given robust test between $\mathcal{U}$ and $\mathcal{V}$ using the black-box robustness-to-privacy transformation. The techniques are from Canonne et al., (2023) and Hopkins et al., (2023).
Lemma C.4.
Suppose $\mathcal{A}$ is a robust test for the mean testing problem (1.2) achieving type-I and type-II errors at most $0.1$ under a contamination level at most $\epsilon$. Then, for $\delta=\frac{1}{\epsilon N}$, there exists a $(0,\delta)$-differentially private algorithm $\widetilde{\mathcal{A}}$ whose error probabilities are also at most $0.1$ for any dataset with size $N$.
@@
Now assume $\widetilde{\mathbf{X}}$ is from $H_{0}$. By the definition of robust testing, with probability at least $0.9$, we can arbitrarily modify $\widetilde{\mathbf{X}}$ and obtain the same result as long as the number of modifications does not exceed $\epsilon N$. This equivalently says that $\mathcal{S}(\widetilde{\mathbf{X}},\mathcal{A})\geq\epsilon N$, which, by the construction of $\widetilde{\mathcal{A}}$, implies that $\widetilde{\mathcal{A}}$ will output $1$ with probability 0 in this case. On the other hand, if $\widetilde{\mathbf{X}}$ is from $H_{1}$, then with probability at least $0.9$, we will have $\mathcal{A}(X)=1$, in which case $\mathcal{S}(X,\mathcal{A})=0$ and consequently $\widetilde{\mathcal{A}}$ will output $1$ with probability $1$. This shows that $\widetilde{\mathcal{A}}$ is a $(0,\delta)$-differential privacy with error probabilities at most $0.1$ under both hypotheses. ∎
We are now prepared to prove the core lemma and Theorem 3.5.
Proof of Lemma C.1.
Suppose we already have a robust test $\mathcal{A}$ for the mean testing problem (1.2) with the fraction of contamination at most $\epsilon$. Then by Lemma C.4, we can contruct a $(0,\delta)$-differential privacy $\widetilde{\mathcal{A}}$ with errors at most $0.1$. Therefore, the constructed algorithm $\widetilde{\mathcal{A}}$ satifies the condition of Lemma C.3. Take $D=\mathbb{E}d_{H}(\mathbf{X},\mathbf{X}^{\prime})$ in Lemma C.3, and by the conclusion of Lemma C.3 we should have $\mathbb{E}d_{H}(\mathbf{X},\mathbf{X}^{\prime})\gtrsim\frac{1}{1/\epsilon N}=\epsilon N$. This necessary condition finishes the proof. ∎
@@
$V:=\left\{v\left\lvert\right.v_{i}=\gamma_{i}\theta_{i},\gamma_{i}\in\left\{-1,1\right\},i\geq 1,i\in I\right\},$
where $\theta$ is the vector from Lemma 3.1 setting $c=\frac{(k_{2}^{\star})^{1/4}\sqrt{\epsilon}}{N^{1/4}}\sigma$, i.e., $\left\lVert\theta\right\rVert_{2}=\frac{(k_{2}^{\star})^{1/4}\sqrt{\epsilon}}{N^{1/4}}\sigma$ and $\left\lVert\theta\right\rVert_{\infty}\leq\frac{\sqrt{\epsilon}}{(k_{2}^{\star})^{1/4}N^{1/4}}\sigma$. We are going to compute $d_{H}(\mathbf{X},\mathbf{X}^{\prime})$ using Lemma B.5. Specifically, we compare (in terms of total variation) a distribution of $k$ exchangeable variables generated from $\mathcal{V}$ to a joint distribution of k1k-1 exchangeable variables generated from $\mathcal{N}\left(v,\sigma^{2}\mathbf{I}\right)$ and the last one from $\mathcal{N}\left(0,\sigma^{2}\mathbf{I}\right)$ for $1\leq k\leq N$. We approach this by decomposing both $\mathbf{X}$ and $\mathbf{X}^{\prime}$ by their sequence of conditional probabilities and constructing the optimal coupling between $X_{k}$ and $X^{\prime}_{k}$ conditional on $X_{1},\dots,X_{k-1}$ using Lemma B.5 such that
@@
where $p(x_{k}|x_{1},\dots,x_{k-1})$ is the density of conditional distribution of the first $k$ coordinates of $X^{\prime}$ and $q(x_{k})$ is the density of standard normal distribution. If we denote the optimal coupling above by $\mathcal{P}_{k}=\mathcal{P}_{k}(X_{k},X^{\prime}_{k}|X_{1},...,X_{k-1})$, it is not hard to verify that
$\widetilde{\mathcal{P}}(\mathbf{X},\mathbf{X}^{\prime}):=\mathcal{P}_{1}(X_{1},X^{\prime}_{1})\cdot\prod\limits_{i=2}^{N}\mathcal{P}_{i}(X_{k},X^{\prime}_{k}|X_{1},\dots,X_{k-1})$
is a valid coupling between $\mathcal{U}$ and $\mathcal{V}$. By the property of the optimal coupling and the definition of total variation distance, we have under $\widetilde{\mathcal{P}}$ that
$\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\sim\widetilde{\mathcal{P}}}\mathbf{1}_{\left\{X_{k}\neq X^{\prime}_{k}\right\}}$ $=\mathbb{E}\text{TV}(p(x_{k}|x_{1},\dots,x_{k-1}),q(x_{k}))$
@@
$\mathbb{E}_{(\mathbf{X},\mathbf{X}^{\prime})\sim\widetilde{\mathcal{P}}}d_{H}(\mathbf{X},\mathbf{X}^{\prime})=\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\sim\widetilde{\mathcal{P}}}\sum\limits_{i=1}^{N}\mathbf{1}_{\left\{X_{i}\neq X^{\prime}_{i}\right\}},$
to evaluate $\mathbb{E}_{(\mathbf{X},\mathbb{X}^{\prime})\sim\widetilde{\mathcal{P}}}d_{H}(\mathbf{X},\mathbf{X}^{\prime})$, it suffices to evaluate
@@
We bound each individual term by the Pinsker’s inequality, where we have
@@
where $(\cdot)_{j}$ denotes the $j$-th coordinate, (i) is from that $\left\lVert v\right\rVert_{2}^{2}$ does not depend on the specific choice of $v\in V$ by the construction of $V$ and (ii) is from an important observation that
@@
where (i) follows from the Jensen’s inequality. Recall that by Lemma 3.1 we have $\left\lVert\theta\right\rVert_{2}=\frac{(k_{2}^{\star})^{1/4}\sqrt{\epsilon}}{N^{1/4}}\sigma$ and $\left\lVert\theta\right\rVert_{\infty}\leq\frac{\sqrt{\epsilon}}{(k_{2}^{\star})^{1/4}N^{1/4}}\sigma$. Now that $\theta_{i}^{4}\leq\theta_{i}^{2}\cdot\left\lVert\theta\right\rVert_{\infty}^{2}$ for all $i\in\left\{1,2,\dots,d\right\}$, we know $\left\lVert\theta\right\rVert_{4}^{2}\leq\left\lVert\theta\right\rVert_{2}\left\lVert\theta\right\rVert_{\infty}\leq\frac{\epsilon}{\sqrt{N}}\sigma^{2}$. By the definition of $V$, we know that $\left\lVert v\right\rVert_{4}^{2}=\left\lVert\theta\right\rVert_{4}^{2}$. Therefore, we have $\frac{\sqrt{N}\left\lVert v\right\rVert_{4}^{2}}{\sigma^{2}}\leq\frac{\sqrt{N}}{\sigma^{2}}\cdot\frac{\epsilon}{\sqrt{N}}\sigma^{2}\leq\epsilon$. By the core lemma, we have
@@
C.4 Proof of Theorem 3.6
In this section, we prove Theorem 3.6. We reach the result by constructing two mixture distributions based on $H_{0}$ and $H_{1}$ which are, in fact, identical. We then evaluate the total variation distance between two truncated binomial distributions, which enables us to construct an adaptive adversary, leading high type-II error. Similar techniques also appear in Prasadan and Neykov, (2025) and Chen et al., (2017).
@@
$q_{1}=\frac{(p_{\mu_{0}}-p_{0})\mathbf{1}_{\left\{p_{\mu_{0}}\geq p_{0}\right\}}}{\text{TV}\left(\mathbb{P}_{0},\mathbb{P}_{\mu_{0}}\right)},\ q_{2}=\frac{(p_{0}-p_{\mu_{0}})\mathbf{1}_{\left\{p_{0}>p_{\mu_{0}}\right\}}}{\text{TV}(\mathbb{P}_{0},\mathbb{P}_{\mu_{0}})}.$ (C.7)
@@
$R_{1}:=(1-\epsilon^{\prime\prime})\mathbb{P}_{0}+\epsilon^{\prime\prime}Q_{1},\ R_{2}:=(1-\epsilon^{\prime\prime})\mathbb{P}_{\mu_{0}}+\epsilon^{\prime\prime}Q_{2}.$ (C.8)
It is not hard to verify that actually $R_{1}=R_{2}$, i.e., they represent the same distribution. Consequently $\text{TV}\left(R_{1}^{\otimes N},R_{2}^{\otimes N}\right)=0$, where $R_{j}^{\otimes N},j\in\left\{1,2\right\}$ is the $N$-fold product of $R_{j}$.
@@
Here $B$ is a binomial random variable with parameters $N$ and $\epsilon^{\prime\prime}$, and (i) follows from the Hoeffding’s inequality.
@@
$0=\text{TV}\left(R_{1}^{\otimes N},R_{2}^{\otimes N}\right)\overset{\text{(ii)}}{\geq}\frac{1}{2}\displaystyle\int_{\mathbb{R}^{d}}\left|S_{1}^{L}-S_{2}^{L}\right|d\mathbf{x}-\frac{1}{2}\displaystyle\int_{\mathbb{R}^{d}}\left|S_{1}^{U}-S_{2}^{U}\right|d\mathbf{x},$
@@
Step IV: conditional argument. Consider a adversary $\mathcal{C}_{B,\epsilon^{\prime\prime}}$ with following strategy. Upon observing the original data $\widetilde{\mathbf{X}}$ and its underlying law (which is either $\mathbb{P}_{0}$ or $\mathbb{P}_{\mu_{0}}$ in our case), the adversary first samples a number $n_{c}$ from the conditional law of $B|B\leq N(\epsilon^{\prime\prime}+1/\sqrt{2N})$, then sample $n_{c}$ replacements from the mixture distribution corresponding to the true underlying law. Specifically, if $\widetilde{\mathbf{X}}$ is from $\mathbb{P}_{0}$, then replacements will be from $Q_{1}$, otherwise the replacements will be from $Q_{2}$. In this way, if we denote $\widetilde{\mathbb{P}}_{0}$ to be the law of the contaminated samples from $\mathbb{P}_{0}$ and $Q_{1}$, and $\widetilde{\mathbb{P}}_{\mu_{0}}$ to be the law of the contaminated samples from $\mathbb{P}_{\mu_{0}}$ and $Q_{2}$, then with minor calculation we have
@@
$\geq 1-\alpha-\text{TV}\left(\widetilde{\mathbb{P}}_{0},\widetilde{\mathbb{P}}_{\mu_{0}}\right)=\varTheta(1).$
@@
Definition C.5 (Consistent subset).
@@
$E_{e}:=\left\{\left\lVert P^{\star}\sum\limits_{i\in T}X_{i}\right\rVert_{2}^{2}-k^{\star}\left|T\right|\sigma^{2}\geq c\left|T\right|^{2}E^{2}(\epsilon,K,N,\sigma^{2})\right\},$ (C.9)